OpenStudy (anonymous):

What type of conic section is given by the equation 4x^2+ 9y^2 = 36? What are its domain and range?

OpenStudy (zzr0ck3r):

well the 4 and 9 tell us its not what?

OpenStudy (anonymous):

No idea.

OpenStudy (zzr0ck3r):

what is this graph x^2 + y^2 = 1?

OpenStudy (anonymous):

not a circle :P

OpenStudy (anonymous):

A circle?

OpenStudy (zzr0ck3r):

its not a circle, its close. what is close to a circle?

OpenStudy (anonymous):

zzr0ck, isnt x^2+y^2=1 a cirlce with center 0,0 and radius 1?

OpenStudy (zzr0ck3r):

yes

OpenStudy (anonymous):

Oh. So its an ellipse.

OpenStudy (zzr0ck3r):

so what would happen if we did 4x^2 + 9x^2 = 1

OpenStudy (zzr0ck3r):

yes @xkylex

OpenStudy (anonymous):

It would be a circle

OpenStudy (anonymous):

\[4x^2+ 9y^2 = 36\]\[{4x^2\over36}+{9y^2\over36}={36\over36}\]\[{x^2\over9}+{y^2\over4}=1\] \[{x^2\over a^2}+{y^2\over b^2}=1~~is~~an~~ellipse\] :)

OpenStudy (anonymous):

zzr0ck that would be an ellipse, but what does the 4 and 9 mean exactly?

OpenStudy (zzr0ck3r):

that is what is stretching it

OpenStudy (anonymous):

@_@, im so confused.

OpenStudy (zzr0ck3r):

a circle is 1*x^2 + 1*y^2 = 1

OpenStudy (anonymous):

so it stretches it how much?

OpenStudy (zzr0ck3r):

when we change the ones it changes

OpenStudy (zzr0ck3r):

that's a hard question, for calculus

OpenStudy (anonymous):

So how would I find the domain and range?

OpenStudy (anonymous):

zzrock if you look at the graph of an ellipse, how can you make the equation from just looking at it??

OpenStudy (zzr0ck3r):

you can get points and solve the equation yummy posted, they always have the same form

OpenStudy (anonymous):

well technically you have to put it into the form of an equation of an ellipse before you graph it...the graph would look like this:

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

The domain would be -3, 3. The range would be -2 2.

OpenStudy (anonymous):

Is that right?

OpenStudy (anonymous):

the domain would be \([-3,3]\) and the range would be \([-2,2]\)

OpenStudy (anonymous):

Thank you everyone! If only I could give more than one medal :c .

OpenStudy (anonymous):

\[a=3\]\[b=2\]\[c=\sqrt{5}\]\[focus:(-\sqrt{5},0)~~and~~(\sqrt{5},0}\]

OpenStudy (zzr0ck3r):

strange question, this is not a function and thus to talk about its domain and range is pointless...

OpenStudy (anonymous):

true, domain and range is pointless, I agree